40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 11 18:02:02 2015, Jenne, Update, LSC, CARM and DARM loops marginal CARMloop_ALSvsREFL_fcav150_fcomp150.pngCARMloop_ALSvsREFL_fcav150_fcomp80.pngDARMloop_ALSvsAS55_fcav4500_nocomp.png
    Reply  Thu Mar 12 03:47:56 2015, Jenne, Update, LSC, CARM and DARM loop adjustments 
Message ID: 11136     Entry time: Thu Mar 12 03:47:56 2015     In reply to: 11133
Author: Jenne 
Type: Update 
Category: LSC 
Subject: CARM and DARM loop adjustments 

No wins tonight.

I've tried playing with the shapes of the loops a little bit (mostly CARM so far), to no avail.  I think I made it to CARM RF-only only one time tonight.  I was able to turn on the REFL11 whitening, although I lost lock while about halfway through the DARM transition. 

I tried making a double integrator instead of a single integrator for CARM_B, since that would allow me to make a complex zero pair which could help win back some phase. I also tried just straight copying FM1 from CARM into CARM_B, so that it could be turned off for the ALS part of the loop, but left on for the REFL part, but that didn't work very well.  Like Rana and I saw last Friday, we really need the REFL signal to have a true integrator, to force the PDH signal toward zero, before we can complete the transition.

I moved the cavity pole compensator's zero back up to 120 Hz, since that was what had worked on Saturday night.  That helped me get farther before running into gain peaking problems at ~50Hz.  This is because, as seen in my simulation earlier tonight, I win back some gain margin by having the pole compensator more closely matched with the pole frequency.

I've been turning off both FM1 and FM2 in CARM and DARM.  I think this is helping a lot, when I can get far enough to do so.  I don't want to turn off the second boost until after I'm about 50/50 on REFL.  (When  I have that much REFL, with the true integrator, the PDH signal sticks to zero).

I tried once turning off the bounce/roll filters for CARM and DARM, rather than the FM1 boost, since the bounce roll filters eat lots of phase, but I got pushed off resonance.  I think not having that focused boost may have made my overall RMS larger, which caused me to randomly jump too far outside of the good PDH range.

Early on in the evening, I turned off the MC2 violin filters in the ETM LSC-SUS filter banks, since I am actuating on only the ETMs tonight. However, I saw a violin mode ring up at ~642, which showed up in POX but not POY.  This was causing up-conversion, beating against the 40-50Hz buzzing from the IR resonance.   The MC2vio1,2 filter covers this frequency (because it's an absurdly wide notch), but the EXYvio1 filter does not.  There seems to be some confusion on the wiki as to what the ETMX violin mode frequency is - it says 631 (638??).  The notch that is in the EXYvio1 filter is for 631 Hz, but this is not correct.  DAYTIME self: Make the MC2 violin filter smaller than 40Hz(!) wide, and move the ETMX notch up to the correct frequency.  For tonight, I just turned back on the MC2 filter, and the mode has rung down.

Idea:  MICH offset, or ETM misalignment, enough to keep the power recycling low-ish, so that the CARM cavity pole doesn't come down too far in frequency?  Daytime brain should think about this.

ELOG V3.1.3-